# supplementary

### v1

@@ Abstract
In high-dimensional linear models, sparsity is often exploited to reduce variability and achieve parsimony. Equi-sparsity, where one assumes that predictors can be aggregated into groups sharing the same effects, is an alternative parsimonious structure that can be more suitable in certain applications. Previous work has clearly demonstrated the benefits of exploiting equi-sparsity in the presence of “rare features” (Yan and Bien, 2021). In this work, we propose a new tree-guided regularization scheme for simultaneous estimation and feature aggregation. Unlike existing methods, our estimator avoids synthetic overparameterization and its detrimental effects. Even though our penalty is applied to hierarchically overlapped groups, we show that its proximal operator can be solved with a one-pass, non-iterative algorithm. Novel techniques are developed to study the finite-sample error bound of this seminorm-induced regularizer under least squares and binomial deviance losses. Theoretically, compared to existing methods, the proposed method offers a faster or equivalent rate depending on the true equi-sparisty structure. Extensive simulation studies verify these findings. Finally, we illustrate the usefulness of the proposed method with an application to a microbiome dataset, where we conduct post-selection inference on the aggregated features’ effects.
@@ 1 Introduction
In microbiome research, modern high‑throughput sequencing experiments quantify the relative abundance of microbes in an environment (e.g., the human gut) by counting the number of rRNA reads sampled that can be assigned to a specific microbe or closely-related group of microbes (measured at the level of an operation taxonomic unit, OTU). For each subject, we obtain a vector $x\in\mathbb{R}^{p}$ with nonnegative entries and unit sum, where $p$ is the number of OTUs in the study. Naturally, the $j$th component of $x$ is the proportion of rRNA reads sampled attributable to the $j$th OTU, which is often interpreted as the relative abundance of that OTU in the subject’s microbiome.
A key analytical goal is to predict a subject-level trait $y$ (e.g., disease status) given $x$ using regression models whose coefficients are then interpreted in terms of associations between OTU abundance and the trait. The high-dimensionality and sparsity of $x$ (typically, many OTUs have zero rRNA reads) motivate researchers to impose sparsity on the coefficients, either by removing features before analysis or by using estimators that shrink coefficients toward zero. Such purely data‑driven feature selection, however, overlooks the biological structure of OTUs and often complicates interpretation. Specifically, OTUs are embedded in a known taxonomic hierarchy reflecting phylogenetic similarity. A natural alternative is thus to aggregate OTUs by taxonomic rank and treat the resulting sums as predictors, representing the relative abundances of higher‑level taxa. The taxonomic tree (e.g., see Figure 7) supports aggregation at various levels (e.g., genus, family, order), reducing dimensionality while yielding biologically meaningful groups, yet the aggregation level is usually chosen ad hoc. This paper introduces a tree‑guided method that adaptively selects aggregation levels based on the data and simultaneously estimates group effects relevant to a given biomedical trait.
@@
Overparameterize-regularize methods typically operate by reparameterizing each regression coefficient as the sum of latent coefficients determined by the tree structure. By imposing sparsity on the latent coefficients, the original coefficients can be naturally grouped to have the same values. Yu and Liu (2016) reparameterize under the guidance of an undirected graph; while Yan and Bien (2021) and Li et al. (2023) designs the latent coefficients based on a tree, lying closer to our problem setting. These methods scale well to high-dimensional data, both computationally and theoretically. However, overparameterization arises several problems, including unfair selection and collinearity, as we will discuss in detail in both Section 3.2 and 7. Moreover, aggregation-driven penalties are naturally seminorm-induced, as the shrinkage center is often a subspace rather than the origin. However, existing works tend to side-step this complication, and only build theory for norm-induced penalties. As a result, no theoretical justification adequately fits the practical usage, especially in cases that adjust for other unpenalized covariates.
In order to address the challenges, we propose a new, direct regularization scheme designed according to the given tree. The proposed method generalizes the aforementioned pairwise fusion penalty to allow larger groups. It avoids the problematic overparameterization of Yan and Bien (2021), and only selects groups that strictly cohere with the tree structure. The benefits are verified in both theoretical and empirical results. We will introduce our method in Section 3, and contrast it against existing methods. In Section 4, we discuss how to compute our estimator efficiently, and illustrate a connection to convex clustering. In Section 5, we present finite-sample consistency results of the proposed estimator under high-dimensional regression and classification settings. We show that the unregularized subspace only impacts prediction error through its dimension in an $O_{P}(n^{-1})$ term. We show that our proposed method can have a better prediction error rate than the method of Yan and Bien (2021) (henceforth, RARE) under certain setting. Moreover, our theory provides insights into the influence of different tree structures on the prediction task and contributes to the literature by providing a novel treatment of seminorm-induced penalties. Simulation studies in Section 6 verify our theoretical findings and show a general advantage over RARE in various settings. Finally, in Section 7, we apply our method to the analysis of microbiome data and illustrate its potential to produce new scientific insights.
Before introducing our method, we use Section 2 to demonstrate the merit feature aggregation.
@@ 2 Benefits of aggregation
[…]
@@
We call the result “condition-free” because it makes no assumptions on the design matrix $X$. The high-probability bound depends only on $\mathop{\rm rank}(X)=r$ and the noise variance $\sigma^{2}$. When $\beta^{*}$ is sparse, dropping zero-effect columns lowers $r$ and improves the bound. Likewise, an improvement can be made under equi-sparsity following a similar mechanism.
@@
[…]
$\frac{1}{\sqrt{n}}\|XH\hat{\beta}_{\rm als}-X\beta^{*}\|_{2}\leq\frac{\sqrt{32\sigma^{2}\log(5)\tilde{r}+32\sigma^{2}\log(r)}}{\sqrt{n}}+\frac{3\|\gamma_{H}\|_{2}}{\sqrt{n}}.$ (2.2)
The result in Proposition 2.2 depicts a trade-off between the bias caused by approximation and the variance controlled by aggregation. The first term characterizes the variance, which is reduced compared to Proposition 2.1 if enough aggregation is conducted so that $\tilde{r}=K<r$. In the high-dimensional regime $n\ll p$ (so $r\leq n$), “sufficient” aggregation means choosing $K<n$, forcing $\tilde{r}=K<r$. The second term is the norm of the approximation error, representing the bias caused by excessive aggregation. So, the aggregated estimator can be better even if the equi-sparsity only holds approximately, especially when the noise variance $\sigma^{2}$ is large. It also aligns with the finding of Yan and Bien (2021) in low-dimensional regression setting with normal noise.
@@ 3.1 Tree-guided convex regularization
Let $\{b_{\ell}\}_{\ell\in[L]}$ be the nodes of a rooted tree (or forest) $\mathcal{T}$, and let $\{b_{\ell}\}_{\ell\in\mathcal{I}}$, for an index set $\mathcal{I}\subset[L]$, be the internal nodes (those with at least one child). The $p$ features correspond to the leaves $\mathcal{L}(\mathcal{T})\subset\{b_{\ell}\}_{\ell\in[L]}$. For any node $b$, write $\mathcal{T}_{b}\subset\mathcal{T}$ for the subtree rooted at $b$ and $\mathcal{L}(\mathcal{T}_{b})\subset\mathcal{L}(\mathcal{T})$ for its leaves, a candidate group of features. We allow $\mathcal{T}$ to be a forest, see Figure 1(c) for an example, whose disconnected components must never be aggregated. Our goal is to discover an “aggregating set” that divides the leaf nodes into groups. We have the following definition from Yan and Bien (Definition 1, 2021).
@@
A node set =$\mathcal{B}=\{b_{1},\dots,b_{k}\}\subset\mathcal{T}$ is an aggregating set if the leaf collections $\bigl{\{}\mathcal{L}(\mathcal{T}_{b_{1}}),\dots,\mathcal{L}(\mathcal{T}_{b_{k}})\bigr{\}}$ form a partition $\mathcal{A}_{\mathcal{B}}=\{\mathcal{A}_{1},\dots,\mathcal{A}_{k}\}$ of $\mathcal{L}(\mathcal{T})$, where $\mathcal{A}_{\ell}\subset[p]:=\{1,\dots,p\}$ indexes the $\ell$th group. If $b_{\ell}$ is itself a leaf, then $\mathcal{L}(\mathcal{T}_{b_{\ell}})=\{b_{\ell}\}$ is a singleton.
Lemma 1 of Yan and Bien (2021) shows that every tree-coherent partition $\mathcal{A}$ admits a unique coarsest aggregating set $\mathcal{B}$: finding a partition is then equivalent to choosing $\mathcal{B}$ and fixing a “resolution” of $\mathcal{T}$. For example, Figure 1 gives two example partitions and represents groups in different colors. In Figure 1(a), the tree-coherent partition $\mathcal{A}=\{\{6\},\{7\},\{1\},\{2,3\},\{4,5\}\}$ yields $\beta=(\beta_{1},\beta_{2},\beta_{2},\beta_{3},\beta_{3},\beta_{4},\beta_{5})^{\top}$ and =$\mathcal{B}=\{b_{6},b_{7},b_{1},b_{8},b_{10}\}$; whereas the partition $\mathcal{A}^{\prime}=\{\{2,3\},\{1,4,5,6,7\}\}$ in Figure 1(b) is incompatible with the tree. We call an incompatible partition tree-incoherent.
Figure 1: (a) A tree-coherent partition with different colors; (b) A tree-incoherent partition; (c) An example of forest structure which can be handled by our method.
@@
$\hat{\beta}\in\operatornamewithlimits{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{g(\beta)+\lambda\,\Omega_{\mathcal{T},w}(\beta)\right\},$ (3.1)
@@
[…]
@@
This penalty sums up the terms for all internal nodes $\{b_{\ell}\}_{\ell\in\mathcal{I}}$ in the tree. For example, in Figure 1(a), there are only four internal nodes $\{b_{11},b_{9},b_{9},b_{10}\}$, corresponding to the groups $\{\mathcal{A}_{{11}}=\{1,2,3,4,5,6,7\},\mathcal{A}_{9}=\{1,2,3\},\mathcal{A}_{{8}}=\{2,3\},$ $\mathcal{A}_{{10}}=\{4,5\}\}$. Then the penalty is $\Omega_{\mathcal{T},w}(\beta)=7^{-1/2}\|D_{7}\beta_{\mathcal{A}_{{11}}}\|_{2}+3^{-1/2}\|D_{3}\beta_{\mathcal{A}_{9}}\|_{2}+2^{-1/2}\|D_{2}\beta_{\mathcal{A}_{{8}}}\|_{2}+2^{-1/2}\|D_{2}\beta_{\mathcal{A}_{{10}}}\|_{2}.$ The penalty is inspired by Molstad and Motwani (2023), where they considered a categorical regression setting, and regularized the coefficients to enforce across-category homogeneity.
@@
In Section 4, we develop an accelerated proximal gradient algorithm (Beck and Teboulle, 2009). The proximal operator of $\Omega_{\mathcal{T},w}$ can be computed with a a one-pass, non-iterative algorithm, so (3.1) can be solved efficiently.
@@ 3.2 Connection to overparameterize-regularize methods
Yan and Bien (2021) induce aggregation through a tree-based reparameterization. They label the $i$th node in the tree with a latent coefficient $\gamma_{i}$ for each $i\in[L]$. For a leaf $j$, its coefficient $\beta_{j}$ is the sum of the latent coefficients of its ancestors, e.g. in Figure 1(a),
$\beta_{1}=\gamma_{8}+\gamma_{9}+\gamma_{1},~~~\beta_{2}=\gamma_{8}+\gamma_{9}+\gamma_{10}+\gamma_{2},~~~\beta_{3}=\gamma_{8}+\gamma_{9}+\gamma_{10}+\gamma_{3},~~\cdots$
This can be summarized as $\beta=A\gamma$ where the expansion matrix $A\in\mathbb{R}^{p\times L}$ has ones in the columns corresponding to each leaf’s ancestors and zeros elsewhere. They estimate $\gamma$ via a LASSO on the reparameterized model
$\hat{\gamma}_{\rm rare}\in\mathop{\rm argmin}_{\gamma\in\mathbb{R}^{L}}\left\{g(A\gamma)+\lambda\sum_{i\in[L]\//\text{roots}}^{L}w_{i}|\gamma_{i}|\right\},~~\hat{\beta}_{\rm rare}=A\hat{\gamma}_{\rm rare},$
where $g$ is the chosen loss (squared error in linear regression) and the $l_{1}$ penalty excludes the root. Because $X\beta=XA\gamma$, each internal node acts as a synthetic predictor equal to the sum of its descendants. Returning to Figure 1(a), with direct selection on synthetic predictors, shrinking $\gamma_{2}$ and $\gamma_{3}$ to zero forces $\beta_{2}=\beta_{3}=\gamma_{8}+\gamma_{9}+\gamma_{10}$; additionally, zeroing $\gamma_{1},\gamma_{10}$ collapses $\beta_{1},\ \beta_{2}$ and $\beta_{3}$ to a common value $\gamma_{8}+\gamma_{9}$, achieving data-driven aggregation. The solution is both neat and can be computed efficiently. Empirically, Bien et al. (2021) report that this RARE strategy outperforms the tree-cut search of Guinot et al. (2018) in both simulations and real data applications.
@@
(i) Tree-incoherent selection. Shrinking arbitrary latent coefficients can aggregate leaves that do not share a subtree (e.g. Fig. 1(b) when only $\gamma_{8},\gamma_{10}$ remain nonzero), undermining tree-based interpretation. This issue is observed in our microbiome data analysis (Section 7). Li et al. (2023) impose a descendant group penalty on the reparameterized parameters to enforce coherence, but at the cost of heavier computation.
(ii) Theory–practice gap. In practice, it is always the case that some latent coefficients are not penalized, e.g., those corresponding to roots in the forest as Figure 1(c). However, both Yan and Bien (2021) and Li et al. (2023) assume every node to be regularized in their theoretical analyses. As such, there is a gap between theory and practice.
@@
In contrast, our method addresses (i)-(iii) by avoiding overparameterization and only penalizing the variability within the candidate aggregating sets implied by the tree. Thus, the selected equi-sparsity structure always coheres with the tree, affording meaningful interpretation. We apply delicate treatment to analyzing the unpenalized subspace and quantify its influence on the prediction error, revealing the theoretical impact of tree structures.

### v2

# A direct approach to tree-guided feature aggregation for high-dimensional regression

@@ Abstract
In high-dimensional linear models, sparsity is often exploited to reduce variability and achieve parsimony. Equi-sparsity, where one assumes that predictors can be aggregated into groups sharing the same effects, is an alternative parsimonious structure that can be more suitable in certain applications. Previous work has clearly demonstrated the benefits of exploiting equi-sparsity in the presence of “rare” features (Yan and Bien, 2021). In this work, we propose a new tree-guided regularization scheme for simultaneous estimation and feature aggregation. Unlike existing methods, our estimator avoids synthetic overparameterization and its detrimental effects. Novel techniques are developed to study the finite-sample error bound of this seminorm-induced regularizer under least squares and binomial deviance losses. Theoretically, compared to existing methods, the proposed method offers a faster or equivalent rate depending on the true equi-sparisty structure. Extensive simulation studies verify these findings. We show that our estimator can be computed very efficiently by exploiting special properties of our penalty. Finally, we illustrate the usefulness of the proposed method with an application to a microbiome dataset, where we conduct post-selection inference on the aggregated features’ effects.
@@ 1 Introduction
In microbiome research, modern high‑throughput sequencing experiments quantify the relative abundance of microbes in an environment (e.g., the human gut) by counting the number of rRNA reads sampled that can be assigned to a specific microbe or closely-related group of microbes (measured at the level of an operational taxonomic unit, OTU). For each subject, these experiments generate a vector $x\in\mathbb{R}^{p}$ with nonnegative entries and unit sum, where $p$ is the number of OTUs in the study. Naturally, the $j$th component of $x$ is the proportion of rRNA reads sampled attributable to the $j$th OTU, which is often interpreted as the relative abundance of that OTU in the subject’s microbiome.
A key analytical goal is to predict a subject-level trait $y$ (e.g., disease status) given $x$ using regression models whose coefficients are then interpreted in terms of associations between OTU abundance and the trait. The high-dimensionality and sparsity of $x$ (typically, many OTUs have zero rRNA reads) motivate researchers to impose sparsity on the coefficients, either by removing features before analysis or by using estimators that shrink coefficients toward zero. Such purely data‑driven feature selection, however, overlooks the biological structure of the OTUs and often complicates interpretation. Specifically, OTUs are embedded in a known taxonomic hierarchy reflecting phylogenetic similarity. A natural alternative is thus to aggregate OTUs by taxonomic rank and treat the resulting sums as predictors, representing the relative abundances of higher‑level taxa. The taxonomic tree (e.g., see Figure 6) supports aggregation at various levels (e.g., genus, family, order), reducing dimensionality while yielding biologically meaningful groups, yet the aggregation level is usually chosen ad hoc. This paper introduces a tree‑guided method that adaptively selects aggregation levels based on the data and simultaneously estimates group effects relevant to a given trait.
@@
Overparameterize-regularize methods typically operate by reparameterizing each regression coefficient as the sum of latent coefficients determined by a known structure. By imposing sparsity on the latent coefficients, the original coefficients can be naturally grouped to have the same values. Yu and Liu (2016) reparameterize under the guidance of an undirected graph; while Yan and Bien (2021) and Li et al. (2023) designs the latent coefficients based on a tree, lying closer to our problem setting. These methods scale well to high-dimensional data, both computationally and theoretically. However, overparameterization arises several problems, including unfair selection and collinearity, as we will discuss in detail in both Section 3.2 and 7. Moreover, aggregation-driven penalties are naturally seminorm-induced, as the shrinkage center is often a subspace rather than the origin. However, existing works tend to side-step this complication, and only build theory for norm-induced penalties. As a result, no theoretical justification adequately fits the practical usage, especially in cases that adjust for other unpenalized covariates.
In order to address the challenges, we propose a new, direct regularization scheme designed according to the given tree. The proposed method generalizes the aforementioned pairwise fusion penalty to allow larger groups. It avoids the problematic overparameterization of Yan and Bien (2021), and only selects groups that strictly cohere with the tree structure. The benefits are verified in both theoretical and empirical results. We will introduce our method in Section 3, and contrast it against existing methods. In Section 4, we discuss how to compute our estimator efficiently. In Section 5, we present finite-sample consistency results of the proposed estimator under high-dimensional regression and classification settings. We show that the unregularized subspace only impacts prediction error through its dimension in an $O_{P}(n^{-1})$ term. We show that our proposed method can have a better prediction error rate than the method of Yan and Bien (2021) (henceforth, RARE) under certain setting. Moreover, our theory provides insights into the influence of different tree structures on the prediction task and contributes to the literature by providing a novel treatment of seminorm-induced penalties. Simulation studies in Section 6 verify our theoretical findings and show a general advantage over RARE in various settings. Finally, in Section 7, we apply our method to the analysis of microbiome data and illustrate its potential to produce new scientific insights. Before introducing our method, we use Section 2 to demonstrate the merit of feature aggregation.
@@ 2 Benefits of aggregation
Assume $y=X\beta^{*}+\epsilon$ with $\epsilon$ having $n$ independent $\sigma$-sub-Gaussian entries. If $\mathop{\rm rank}(X)=r$, then with probability at least 12/r1-2/r,
@@
We call the preceding result “condition-free” as it makes no assumptions on the design matrix $X$. The high-probability bound depends only on $\mathop{\rm rank}(X)=r$ and the noise variance $\sigma^{2}$. When $\beta^{*}$ is sparse, dropping zero-effect columns lowers $r$ and improves the bound. Likewise, an improvement can be made under equi-sparsity following a similar mechanism.
@@
Under the assumptions of Proposition 2.1, we have $\mathop{\rm rank}(\tilde{X})=\tilde{r}\leq\min(r,K)$. Then, with probability at least 12/r1-2/r,
$\vskip-10.0pt\frac{1}{\sqrt{n}}\|XH\hat{\beta}_{\rm als}-X\beta^{*}\|_{2}\leq\frac{\sqrt{32\sigma^{2}\log(5)\tilde{r}+32\sigma^{2}\log(r)}}{\sqrt{n}}+\frac{3\|\gamma_{H}\|_{2}}{\sqrt{n}}.$ (2.2)
The result in Proposition 2.2 illustrates a trade-off between the variance controlled by aggregation and the bias caused by excessive aggregation. Specifically, the first term characterizes the variance, which is reduced compared to Proposition 2.1 if enough aggregation is conducted so that $\tilde{r}=K<r$. In the high-dimensional regime $n\ll p$ (so $r\leq n$), “sufficient” aggregation means choosing $K<n$, forcing $\tilde{r}=K<r$. The second term is the norm of the approximation error, representing the bias caused by excessive aggregation. Evidently, the aggregated estimator can be better even if equi-sparsity only holds approximately, especially when the noise variance $\sigma^{2}$ is large. This coheres with the findings of Yan and Bien (2021) in low-dimensional regression setting with normal noise.
@@ 3.1 Tree-guided convex regularization
Let $\{b_{\ell}\}_{\ell\in[L]}$ be all the nodes of a rooted tree (or forest) $\mathcal{T}$, and let $\{b_{\ell}\}_{\ell\in\mathcal{I}}$, for an index set $\mathcal{I}\subset[L]$, be the internal nodes (those with at least one child). The $p$ features correspond to the leaves $\mathcal{L}(\mathcal{T})\subset\{b_{\ell}\}_{\ell\in[L]}$. For any node $b$, write $\mathcal{T}_{b}\subset\mathcal{T}$ for the subtree rooted at $b$ and $\mathcal{L}(\mathcal{T}_{b})\subset\mathcal{L}(\mathcal{T})$ for its leaves, a candidate group of features. We allow $\mathcal{T}$ to be a forest (see Figure 1(a), for example) whose disconnected components must never be aggregated. Our goal is to discover an “aggregating set” that divides the leaf nodes into groups. We have the following definition from Yan and Bien (Definition 1, 2021).
@@
[…]
Lemma 1 of Yan and Bien (2021) shows that every tree-coherent partition $\mathcal{A}$ admits a unique coarsest aggregating set $\mathcal{B}$: finding a partition is then equivalent to choosing $\mathcal{B}$ and fixing a “resolution” of $\mathcal{T}$. For example, Figure 1 gives two example partitions and represents groups in different colors. In Figure 1(b), the tree-coherent partition $\mathcal{A}=\{\{6\},\{7\},\{1\},\{2,3\},\{4,5\}\}$ yields $\beta=(\beta_{1},\beta_{2},\beta_{2},\beta_{3},\beta_{3},\beta_{4},\beta_{5})^{\top}$ and $\mathcal{B}=\{b_{6},b_{7},b_{1},b_{8},b_{10}\}$; whereas the partition $\mathcal{A}^{\prime}=\{\{2,3\},\{1,4,5,6,7\}\}$ in Figure 1(c) is incompatible with the tree. We call an incompatible partition tree-incoherent.
Figure 1: (a) An example of forest structure which can be handled by our method; (b) A tree-coherent partition where leaf nodes of the same color are aggregated; (c) A tree-incoherent partition where leaf nodes of the same color are aggregated.
@@
$\vskip-10.0pt\hat{\beta}\in\operatornamewithlimits{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\{g(\beta)+\lambda\,\Omega_{\mathcal{T},w}(\beta)\right\},$ (3.1)
@@
$\vskip-10.0pt\Omega_{\mathcal{T},w}(\beta)\;=\;\sum_{\ell\in\mathcal{I}}w_{\ell}\min_{c\in\mathbb{R}}\bigl\|\beta_{\mathcal{A}_{\ell}}-c\,\mathbf{1}_{a_{\ell}}\bigr\|_{2}\;=\;\sum_{\ell\in\mathcal{I}}w_{\ell}\bigl\|D_{a_{\ell}}\beta_{\mathcal{A}_{\ell}}\bigr\|_{2},$ (3.2)
@@
This penalty sums up the terms for all internal nodes $\{b_{\ell}\}_{\ell\in\mathcal{I}}$ in the tree. For example, in Figure 1(b), there are only four internal nodes $\{b_{11},b_{9},b_{8},b_{10}\}$, corresponding to the groups $\{\mathcal{A}_{{11}}=\{1,2,3,4,5,6,7\},\mathcal{A}_{9}=\{1,2,3\},\mathcal{A}_{{8}}=\{2,3\},$ $\mathcal{A}_{{10}}=\{4,5\}\}$. Then the penalty is $\Omega_{\mathcal{T},w}(\beta)=7^{-1/2}\|D_{7}\beta_{\mathcal{A}_{{11}}}\|_{2}+3^{-1/2}\|D_{3}\beta_{\mathcal{A}_{9}}\|_{2}+2^{-1/2}\|D_{2}\beta_{\mathcal{A}_{{8}}}\|_{2}+2^{-1/2}\|D_{2}\beta_{\mathcal{A}_{{10}}}\|_{2}.$ The penalty is inspired by Molstad and Motwani (2023), where they considered a categorical regression setting, and regularized the coefficients to enforce across-category homogeneity.
@@
In Section 4, we develop an accelerated proximal gradient algorithm (Beck and Teboulle, 2009). The proximal operator of $\Omega_{\mathcal{T},w}$ can be computed with a one-pass, non-iterative algorithm, so (3.1) can be solved efficiently.
@@ 3.2 Connection to overparameterize-regularize methods
Yan and Bien (2021) induce aggregation through a tree-based reparameterization. They label the $i$th node in the tree with a latent coefficient $\gamma_{i}$ for each $i\in[L]$. For a leaf $j$, its coefficient $\beta_{j}$ is the sum of the latent coefficients of its ancestors, e.g. in Figure 1(b), $\beta_{1}=\gamma_{8}+\gamma_{9}+\gamma_{1},\beta_{2}=\gamma_{8}+\gamma_{9}+\gamma_{10}+\gamma_{2},\beta_{3}=\gamma_{8}+\gamma_{9}+\gamma_{10}+\gamma_{3},$ and so on. This can be summarized as $\beta=A\gamma$ where the expansion matrix $A\in\mathbb{R}^{p\times L}$ has ones in the columns corresponding to each leaf’s ancestors and zeros elsewhere. They estimate $\gamma$ via a LASSO on the reparameterized model
$\vskip-10.0pt\hat{\gamma}_{\rm rare}\in\mathop{\rm argmin}_{\gamma\in\mathbb{R}^{L}}\left\{g(A\gamma)+\lambda\sum_{i\in[L]\//\text{roots}}^{L}w_{i}|\gamma_{i}|\right\},~~\hat{\beta}_{\rm rare}=A\hat{\gamma}_{\rm rare},$
where $g$ is the chosen loss (squared error in linear regression) and the $l_{1}$ penalty excludes the root. Because $X\beta=XA\gamma$, each internal node acts as a synthetic predictor equal to the sum of its descendants. Returning to Figure 1(b), with direct selection on synthetic predictors, shrinking $\gamma_{2}$ and $\gamma_{3}$ to zero forces $\beta_{2}=\beta_{3}=\gamma_{8}+\gamma_{9}+\gamma_{10}$; additionally, zeroing $\gamma_{1},\gamma_{10}$ collapses $\beta_{1},\ \beta_{2}$ and $\beta_{3}$ to a common value $\gamma_{8}+\gamma_{9}$, achieving data-driven aggregation. The solution is both neat and can be computed efficiently. Empirically, Bien et al. (2021) report that this RARE strategy outperforms the tree-cut search of Guinot et al. (2018) in both simulations and real data applications.
@@
(i) Tree-incoherent selection. Shrinking arbitrary latent coefficients can aggregate leaves that do not share a subtree (e.g. Fig. 1(c) when only $\gamma_{8},\gamma_{10}$ remain nonzero), undermining tree-based interpretation. This issue is observed in our microbiome data analysis (Section 7). Li et al. (2023) impose a descendant group penalty on the reparameterized parameters to enforce coherence, but at the cost of heavier computation.
(ii) Theory–practice gap. In practice, it is always the case that some latent coefficients are not penalized, e.g., those corresponding to roots in the forest as Figure 1(a). However, both Yan and Bien (2021) and Li et al. (2023) assume every node to be regularized in their theoretical analyses. As such, there is a gap between theory and practice.
@@
In contrast, our method addresses (i)-(iii) by avoiding overparameterization and only penalizing the variability within the candidate aggregating sets implied by the tree. Thus, the selected equi-sparsity structure always coheres with the tree, affording meaningful interpretation. We carefully analyzing the unpenalized subspace and quantify its influence on the prediction error, revealing the theoretical impact of tree structures.